OpenStudy (anonymous):

PLEASE HELP d = dCosB/CosA + yCosB I need to prove that it equals yCosACosB/CosA - CosB Just give me help on how to get it like this, step by step. Any help is greatly appreciated :)

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (jdoe0001):

so we can see what you mean

OpenStudy (anonymous):

ok wait a sec

OpenStudy (anonymous):

@jdoe0001 thank you for having a look :)

OpenStudy (jdoe0001):

hmmm I'm assuming there's a context to it was wondering what "d" and "y" were =)

OpenStudy (anonymous):

d is the distance, don't know what y is. would you like me to give you the original question?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

the question

OpenStudy (anonymous):

working so far

OpenStudy (anonymous):

does that help? @jdoe0001

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@whpalmer4 @jdoe0001 are you guys having any luck?

OpenStudy (jdoe0001):

hehe, not yet

OpenStudy (whpalmer4):

Okay, it's not so bad. \[\cos (\alpha )=\frac{d}{x}\]\[x = \frac{d}{\cos(\alpha)}\]\[d = x\cos(\alpha)\] \[\cos (\beta )=\frac{d}{x+y}\]\[d = (x+y)\cos(\beta)\] \[x\cos(\alpha) = (x+y)\cos(\beta)\]\[x\cos(\alpha)=x\cos(\beta) + y\cos(\beta)\]\[x(\cos(\alpha)-\cos(\beta)) = y\cos(\beta)\] Now divide both sides by the difference of cosines term from the left, and substitute what we got for \(x\) in the second step. Then multiply by \(cos(\alpha)\) and Bob's yer uncle...

OpenStudy (anonymous):

@whpalmer4 you are a lifesaver thank you so so much!! could i just show you a photo of the working out ot make sure its correct?? :) and my teacher substituted x=d/cosA into it, but you substiuted d=xcosA, would that mke any difference?

OpenStudy (whpalmer4):

\[x = \frac{d}{\cos(\alpha)}\]\[x*\cos(\alpha) = \frac{d}{\cancel{\cos(\alpha)}}*\cancel{\cos(\alpha)}=d\]

OpenStudy (whpalmer4):

sure, show me a picture

OpenStudy (anonymous):

ok, um i just don't get when you wrote xcosA = xcosB +ycosB xcosA -cosB = ycosB

OpenStudy (jdoe0001):

\(\bf cos(\alpha)=\cfrac{d}{x}\implies xcos(\alpha)={\color{red}{ d}}\qquad {\color{blue}{ x}}=\cfrac{d}{cos(\alpha)} \\ \quad \\ cos(\beta)=\cfrac{d}{x+y}\implies xcos(\beta)+ycos(\beta)={\color{red}{ d}} \\ \quad \\ xcos(\alpha)=xcos(\beta)+ycos(\beta)\implies xcos(\alpha)-xcos(\beta)=ycos(\beta) \\ \quad \\ x[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \implies d=\cfrac{ycos(\beta)}{cos(\alpha)-cos(\beta)}\)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Look at the penultimate line of @jdoe0001 's writeup to see what I did in the "mystery step" :-)

OpenStudy (whpalmer4):

@rhiannon1 you dropped the \(x\) off the \(x\cos(\beta)\) term when you moved it across the = sign.

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ \implies \cfrac{d[cos(\alpha)-cos(\beta)]}{cos(\alpha)}=ycos(\beta) \\ \quad \\ \implies d=\cfrac{ycos(\beta)}{cos(\alpha)-cos(\beta)}\)

OpenStudy (anonymous):

yeah, i didnt get it but now i see that you have to take the xcosB across the equals sign then factorise, right?

OpenStudy (jdoe0001):

common factor, yes

OpenStudy (anonymous):

@jdoe0001 i dont get what you wrote in the message before the most recent one

OpenStudy (jdoe0001):

ohh just an expanded version of the last line in case you didn't see the simplification :)

OpenStudy (whpalmer4):

yes, that's what I combined into 1 step, moving the term and factoring out the \(x\).

OpenStudy (anonymous):

but its different ot the answer @jdoe0001

OpenStudy (anonymous):

to*

OpenStudy (jdoe0001):

hmmmm

OpenStudy (whpalmer4):

he forgot to put in the \(\cos(\alpha)\) that he multiplied both sides by \[\bf {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ \implies \cfrac{d[cos(\alpha)-cos(\beta)]}{cos(\alpha)}=ycos(\beta) \\ \quad \\ \implies d=\cfrac{y\cos(\alpha)cos(\beta)}{cos(\alpha)-cos(\beta)}\]

OpenStudy (jdoe0001):

ohh shoot... ahemm yea =)

OpenStudy (whpalmer4):

I'm still waiting for the instantaneous thought->LaTeX converter to arrive :-)

OpenStudy (jdoe0001):

\(\bf \bf cos(\alpha)=\cfrac{d}{x}\implies xcos(\alpha)={\color{red}{ d}}\qquad {\color{blue}{ x}}=\cfrac{d}{cos(\alpha)} \\ \quad \\ cos(\beta)=\cfrac{d}{x+y}\implies xcos(\beta)+ycos(\beta)={\color{red}{ d}} \\ \quad \\ xcos(\alpha)=xcos(\beta)+ycos(\beta)\implies xcos(\alpha)-xcos(\beta)=ycos(\beta) \\ \quad \\ x[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \implies d=\cfrac{y{\color{blue}{ cos(\alpha)}}cos(\beta)}{cos(\alpha)-cos(\beta)}\)

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

rhiannon1 just had a typo =)

OpenStudy (anonymous):

oh ok, but how did you get the denominator like that?

OpenStudy (jdoe0001):

\(\bf cos(\alpha)=\cfrac{d}{x}\implies xcos(\alpha)={\color{red}{ d}}\qquad {\color{blue}{ x}}=\cfrac{d}{cos(\alpha)} \\ \quad \\ ----------------------\\ {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ \implies \cfrac{{\color{blue}{ d}}[cos(\alpha)-cos(\beta)]}{{\color{blue}{ cos(\alpha)}}}=ycos(\beta) \\ \quad \\ \implies {\color{blue}{ d}}=\cfrac{y{\color{blue}{ cos(\alpha)}}cos(\beta)}{cos(\alpha)-cos(\beta)} \) see it?

OpenStudy (anonymous):

not really :(

OpenStudy (jdoe0001):

\(\bf cos(\alpha)=\cfrac{d}{x}\implies xcos(\alpha)={\color{red}{ d}}\qquad {\color{blue}{ x}}=\cfrac{d}{cos(\alpha)} \\ \quad \\ ----------------------\\ {\color{blue}{ \cfrac{d}{cos(\alpha)}}}[cos(\alpha)-cos(\beta)]=ycos(\beta) \\ \quad \\ \implies \cfrac{{\color{blue}{ d}}[cos(\alpha)-cos(\beta)]}{{\color{blue}{ cos(\alpha)}}}=ycos(\beta) \\ \quad \\ \implies {\color{blue}{ d}}[cos(\alpha)-cos(\beta)]=y{\color{blue}{ cos(\alpha)}}cos(\beta) \\ \quad \\ \implies {\color{blue}{ d}}=\cfrac{y{\color{blue}{ cos(\alpha)}}cos(\beta)}{cos(\alpha)-cos(\beta)}\)

OpenStudy (jdoe0001):

up to what part you understand it?

OpenStudy (anonymous):

i think i get it now that you did every step. so you time by cosa then divide by the part in brackets?

OpenStudy (jdoe0001):

yes, more or less you'd cross-multiply to get the \(cos(\alpha)\) over then cross-multiply to get the bracket group down

OpenStudy (anonymous):

@jdoe0001 @whpalmer4 do you think this is year 10 work? just wondering.

OpenStudy (jdoe0001):

10th grade you mean?

OpenStudy (anonymous):

yep

OpenStudy (jdoe0001):

hmmmm might be ambiguous depends on how much you've covered trig and what has been over in the book but maybe or maybe not

OpenStudy (anonymous):

ok, don't know what ambiguous means, but ive heard of it.

OpenStudy (jdoe0001):

ambiguous, it can go either way

OpenStudy (anonymous):

when you started the working out, from the d/cosa (cosa - cosB) = ycosB i dont see how you did that (this is my last question) i mean where did you get it from?

OpenStudy (anonymous):

oh wait, did you sub in the x = d/cosa into x(cosa-cosB) = ycosB ??@jdoe0001

OpenStudy (whpalmer4):

yes

OpenStudy (anonymous):

oh ok, thank you guys so much! i know who to ask for math problems haha thank you for your time i understand a lot better thankyou!! :)

OpenStudy (whpalmer4):

\[\cos (\alpha )=\frac{d}{x}\]\[x = \frac{d}{\cos(\alpha)}\]\[d = x\cos(\alpha)\]\[\cos (\beta )=\frac{d}{x+y}\]\[d = (x+y)\cos(\beta)\] \[x\cos(\alpha) = (x+y)\cos(\beta)\]\[x\cos(\alpha)=x\cos(\beta) + y\cos(\beta)\]\[x\cos(\alpha)-x\cos(\beta) = y\cos(\beta)\]\[x(\cos(\alpha)-\cos(\beta)) = y\cos(\beta)\]Now substitute for \(x\)\[\frac{d}{\cos(\alpha)}*(\cos(\alpha)-\cos(\beta)) = y\cos(\beta)\] \[\frac{d}{\cos(\alpha)}= \frac{y\cos(\beta)}{\cos(\alpha)-\cos(\beta)}\]Multiply both sides by \(\cos(\alpha)\)\[d = \frac{y\cos(\alpha)\cos(\beta)}{\cos(\alpha)-\cos(\beta)}\]

OpenStudy (whpalmer4):

I think this is maybe on the slightly harder end of a high school trigonometry class, but not too much. We didn't have to do any real heavy lifting here with multiple trig identities or anything like that.

OpenStudy (whpalmer4):

Mostly, it was just an exercise in algebra after you drew the picture, right?

OpenStudy (anonymous):

yeah @whpalmer4

OpenStudy (whpalmer4):

A tactic I occasionally find helpful (thought about doing it here, but didn't want to have to explain it while you were trying to understand the problem) is to replace all the trig functions with letters. We aren't doing any trig identity work here, so it isn't crucial that you be able to recognize them as such, and often the algebra becomes much simpler to see (and write/typeset!) I also do this with algebra where I have a bunch of square roots or other radical signs...

OpenStudy (whpalmer4):

Then you undo the substitution after you've done all of the shuffling around and take another look to see if there is anything more you can do.